# On non-existence of bifurcations in one-dimensional Bratu equation

### v1

@@ 1 Introduction
The Bratu (Liouiville-Bratu-Gelfand) equation serves as a model in several applications, including the modeling of thermal ignition [Frank-Kamenetskii] and chemical diffusion processes [Frank-Kamenetskii, JACOBSEN2002283]. It has been used in the theory of chemical reactors [Gavalas1968] and as a model for the temperature distribution of an isothermal gas sphere in gravitational equilibrium [Chandrashekhar]. The Bratu equation is employed as a prototypical example to test the capability of a numerical method to compute multiple solutions to nonlinear parameterized boundary value problems, see [Farrell2015] for example.
@@
satisfying Dirichlet boundary conditions $u(x=\pm 1)=0$. It is well known that the numerically computed bifurcation diagram for (1) using the finite difference (FD) method gives rise to spurious bifurcations [govaerts2000numerical, korman2003accurate]. Spurious solutions for the difference equations arising from the discretization are not limited to second order differential equations (ODEs) of the form (1). Such solutions were obtained by Domokos and Holmes [domokos2003nonlinear] for the ODE that governs beam buckling. Allgower [ALLGOWER19751] has shown that $u^{\prime\prime}+\lambda u^{k}=0$, $u(0)=u(\pi)=0$ may admit what they refer to as ‘numerically irrelevant’ solutions (NIS) when the FD approximation is used. More generally, Peitgen [peitgen1980nonlinear] and Beyn and Lorenz [beyn1982spurious] reported the existence of NIS in FD solutions for various boundary value problems of the form $u^{\prime\prime}+\lambda f(u)=0$, where $f$ is an asymptotically linear or superlinear continuous function having several zeros. In particular, they showed that a plethora of spurious solutions are obtained when $f(u)=e^{u}$. Kuiper [kuiper1988spurious] investigated the discretized version of the FD of (1) and proved that spurious solutions bifurcating from the branch of symmetric solutions are antisymmetric about $x=0$. Moreover, it was shown that these bifurcations occur only when an odd number of elements are used in discretization.
The above discussion regarding the occurrence of spurious bifurcations in the discretized version of the Bratu equation serves as a primary motivation to study the bifurcations of (1). The objectives of this work are two fold: (1) We report on spurious bifurcations in finite element (FE) approximation to the Bratu equation, which has not been previously reported in the literature to the best of our knowledge. (2) Using the tools of analytical bifurcation theory, we present a direct proof showing that there are no bifurcations on the non-trivial branch of solutions to the continuous one-dimensional Bratu equation. The result is significant given the rarity of problems in which bifurcations along a non-trivial global solution branch can be ruled out.
@@ 2 Numerical results
The objective of this section is to show the existence of spurious bifurcation points on the numerically computed bifurcation diagram for (1). We present these results using the FE method in addition to the FD method reported in the literature (see [govaerts2000numerical]). For each of the two methods, we use a discretization of the domain $x\in[-1,1]$ with $N$ number of elements such that the grid is uniformly spaced with grid size $h=2/N$ and grid points placed symmetrically about $x=0$. We then employ Newton’s method coupled with the numerical continuation technique [keller1987lectures] to solve the resulting parametrized nonlinear algebraic equations.
@@
Clearly, $(u(x),\lambda)\equiv(0,0)$ corresponds to the trivial solution of (1). The non-trivial solutions of (1), denoted by $(u_{0}(x),\lambda_{0})$, computed using the FE and FD methods are plotted in bifurcation diagrams shown in Fig. 1(a) and Fig. 1(b) respectively, with the response $u^{*}:=u(x=0)$ on the vertical axis and the parameter $\lambda$ on the horizontal axis. At every solution point along the bifurcation path, we monitor the eigenvalues of the tangent stiffness matrix. The occurrence of a zero eigenvalue on the path indicates a critical point (a limit point or a bifurcation point). Starting with $(u_{0},\lambda_{0})=(0,0)$, a zero eigenvalue is first observed at the limit point. Upon continuing along the upper branch, we encounter a second critical point corresponding to a spurious bifurcation point, shown with a red circle for both FE and FD methods. These bifurcation points are observed only when an odd number ($N$) of elements are used in the discretization. The results corresponding to $N=3,5,\text{and}\ 7$ are shown in Figures 1(a) and 1(b). Observe that the bifurcation point moves upwards along increasing $u_{0}^{*}$ as the value of $N$ increases. The exact location of bifurcation points for a few selected number of elements up to $N=101$ using FE and FD are given in Tables 1 and 2 respectively. We conclude that, for both the discretization schemes, the bifurcation point does not converge with increasing numerical refinement, it rather differs significantly for larger values of $N$. This is one of the features of spurious bifurcations, highlighted using FD results in [peitgen1980nonlinear].
@@
Another signature of spurious bifurcations is the lack of reflection symmetry of the eigenvector corresponding to the zero eigenvalue [govaerts2000numerical]. This can be seen in Figure 1(c) which shows the eigenvector $w(x)$, having a characteristic “saw-tooth” pattern antisymmetric about $x=0$, for $N=101$ elements using FE. Similar antisymmetric eigenvectors were obtained for other $N$ values considered in this work, for both FE and FD implementation schemes. Note, however, that the numerical solutions $u(x)$ along the non-trivial solution path are symmetric about $x=0$ for all parameter values.
@@ 3 Analysis of critical points
In this section, we present an analytical proof demonstrating the non-existence of bifurcation points by directly working with the continuous Bratu equation. In particular, we suggest a change of variable that renders the variable coefficient ODE, resulting from the linearization of (1) about the non-trivial solution, amenable for analysis. To this end, we begin by recalling the well-known non-trivial solutions of (1) [liouville1853equation]. They can be written in terms of $u_{0}^{*}\in(0,\infty)$ as,
@@
To proceed further, we introduce a parameter $\alpha$ such that
@@
and rewrite (2) in parameterized form as
@@
Satisfaction of (14) requires that $c_{1}\tanh{\alpha}=0$. As $\alpha>0$, we get $c_{1}=0$. For nontrivial solutions, the boundary conditions on $w(x)$ require that
@@
Therefore, the kernel of (10) is spanned by
@@
Remark. Note that, maximizing $\lambda_{0}$ given in (4) with respect to $\alpha$ gives the condition (15) for the limit point alone. However, we have shown that (15) is indeed a necessary condition for the existence of critical points, i.e. bifurcation points and limit points in general.
@@
then there exists a unique continuously differentiable curve $((u_{0}(s),\lambda_{0}(s))|s\in(-\delta,\delta)),(u_{0}(0),\lambda_{0}(0))=(\bar{u}_{0},\bar{\lambda}_{0})$ such that $F(u_{0}(s),\lambda_{0}(s))\equiv 0$ [kielhofer2011].
@@ 4 Concluding remarks
The bifurcation diagrams in Section 2 were obtained considering the direct discretization of the nonlinear Bratu equation given in (1). An alternate approach to computing the bifurcation points involves tracking the non-trivial solutions of the discretized linear differential equations (10) or (11) for $\alpha\in(0,\infty)$. While the FD discretized (10) results in non-trivial solutions corresponding to spurious bifurcations, the FD discretized Legendre equation (11) does not have non-trivial solutions for any $\alpha$ other than the limit point. Indeed, a change of variable eliminated their occurrence in the discrete equations, consistent with the analytical results for the continuous case.
We highlight that, even with a choice of discretization that respects the reflection symmetry required by (1), we get spurious bifurcations that yield antisymmetric eigenmodes, although the numerical solutions $u(x)$ in these cases are all symmetry-preserving. In fact, the numerical bifurcation paths with coarser FE discretizations agree well with (2) except near the limit point. Through this work, we bring to attention that while numerical approaches are routinely used for parametrized nonlinear problems, a systematic verification of the numerical bifurcations cannot be discounted.

### v2

@@ 1 Introduction
The Bratu (Liouiville-Bratu-Gelfand) equation serves as a model in several applications, including the modeling of thermal ignition [6] and chemical diffusion processes [6, 9]. It has been used in the theory of chemical reactors [7] and as a model for the temperature distribution of an isothermal gas sphere in gravitational equilibrium [3]. The Bratu equation is employed as a prototypical example to test the capability of a numerical method to compute multiple solutions to nonlinear parameterized boundary value problems, see [5] for example.
@@
satisfying Dirichlet boundary conditions $u(x=\pm 1)=0$. It is well known that the numerically computed bifurcation diagram for (1) using the finite difference (FD) method gives rise to spurious bifurcations [8, 12]. Spurious solutions for the difference equations arising from the discretization are not limited to second order differential equations (ODEs) of the form (1). Such solutions were obtained by Domokos and Holmes [4] for the ODE that governs beam buckling. Allgower [1] has shown that $u^{\prime\prime}+\lambda u^{k}=0$, $u(0)=u(\pi)=0$ may admit what they refer to as ‘numerically irrelevant’ solutions (NIS) when the FD approximation is used. More generally, Peitgen [15] and Beyn and Lorenz [2] reported the existence of NIS in FD solutions for various boundary value problems of the form $u^{\prime\prime}+\lambda f(u)=0$, where $f$ is an asymptotically linear or superlinear continuous function having several zeros. In particular, they showed that a plethora of spurious solutions are obtained when $f(u)=e^{u}$. Kuiper [13] investigated the discretized version of the FD of (1) and proved that spurious solutions bifurcating from the branch of symmetric solutions are antisymmetric about $x=0$. Moreover, it was shown that the bifurcations occur only when an odd number of elements are used in discretization.
The above discussion regarding the occurrence of spurious bifurcations in the discretized version of the Bratu equation serves as a primary motivation to study the bifurcations of (1). The objectives of this work are two fold: (1) We report on spurious bifurcations in finite element (FE) approximation to the Bratu equation, which has not been previously reported in the literature to the best of our knowledge. (2) Using the tools of analytical bifurcation theory, we present a direct proof showing that there are no bifurcations on the non-trivial branch of solutions to the continuous one-dimensional Bratu equation. We emphasize that by bifurcation we necessarily mean branching of solutions[16, 17]. Therefore, our result is significant given the rarity of problems in which bifurcations along a non-trivial global solution branch can be ruled out.
@@ 2 Numerical results
The objective of this section is to show the existence of spurious bifurcation points on the numerically computed bifurcation diagram for (1). We present these results using the FE method in addition to the FD method reported in the literature (see [8]). For each of the two methods, we use a discretization of the domain $x\in[-1,1]$ with $N$ number of elements such that the grid is uniformly spaced with grid size $h=2/N$ and grid points placed symmetrically about $x=0$. We then employ Newton’s method coupled with the numerical continuation technique [10] to solve the resulting parametrized nonlinear algebraic equations.
@@
Clearly, $(u(x),\lambda)\equiv(0,0)$ corresponds to the trivial solution of (1). The non-trivial solutions of (1), denoted by $(u_{0}(x),\lambda_{0})$, computed using the FE and FD methods are plotted in bifurcation diagrams shown in Fig. 1(a) and Fig. 1(b) respectively, with the response $u^{*}:=u(x=0)$ on the vertical axis and the parameter $\lambda$ on the horizontal axis. At every solution point along the bifurcation path, we monitor the eigenvalues of the tangent stiffness matrix. The occurrence of a zero eigenvalue on the path indicates a critical point (a limit point or a bifurcation point). Starting with $(u_{0},\lambda_{0})=(0,0)$, a zero eigenvalue is first observed at the limit point. Upon continuing along the upper branch, we encounter a second critical point corresponding to a spurious bifurcation point, shown with a red circle for both FE and FD methods. These bifurcation points are observed only when an odd number ($N$) of elements are used in the discretization. The results corresponding to $N=3,5,\text{and}\ 7$ are shown in Figures 1(a) and 1(b). Observe that the bifurcation point moves upwards along increasing $u_{0}^{*}$ as the value of $N$ increases. The exact location of bifurcation points for a few selected number of elements up to $N=101$ using FE and FD are given in Tables 1 and 2 respectively. We conclude that, for both the discretization schemes, the bifurcation point does not converge with increasing numerical refinement, it rather differs significantly for larger values of $N$. This is one of the features of spurious bifurcations, highlighted using FD results in [15].
@@
Another signature of spurious bifurcations is the lack of reflection symmetry of the eigenvector corresponding to the zero eigenvalue [8]. This can be seen in Figure 1(c) which shows the eigenvector $w(x)$, having a characteristic “saw-tooth” pattern antisymmetric about $x=0$, for $N=101$ elements using FE. Similar antisymmetric eigenvectors were obtained for other $N$ values considered in this work, for both FE and FD implementation schemes. Note however that, the numerical solutions $u(x)$ along the non-trivial solution path are symmetric about $x=0$ for all parameter values.
@@ 3 Analysis of critical points
In this section, we present an analytical proof demonstrating the non-existence of bifurcation points by directly working with the continuous Bratu equation. In particular, we suggest a change of variable that renders the variable coefficient ODE, resulting from the linearization of (1) about the non-trivial solution, amenable for analysis. To this end, we begin by recalling the well-known non-trivial solutions of (1) [14]. They can be written in terms of $u_{0}^{*}\in(0,\infty)$ as,
@@
To proceed further, we introduce a parameter $\alpha$ such that,
@@
and rewrite (2) in parameterized form as,
@@
Satisfaction of (14) requires that $c_{1}\tanh{\alpha}=0$. As $\alpha>0$, we get $c_{1}=0$. For nontrivial solutions, the boundary conditions on $w(x)$ require that,
@@
Therefore, the kernel of (10) is spanned by,
@@
Remark. Note that, maximizing $\lambda_{0}$ given in (4) with respect to $\alpha$ gives the condition (15) for the limit (turning) point alone. However, we have shown that (15) is indeed a necessary condition for the existence of all critical points, i.e. bifurcation points and limit points in general. We now proceed to show that the critical point obtained using 15 is a limit point, ruling out any bifurcations (branching) on the primary path of solutions 4.
@@
then there exists a unique continuously differentiable curve $((u_{0}(s),\lambda_{0}(s))|s\in(-\delta,\delta)),(u_{0}(0),\lambda_{0}(0))=(\bar{u}_{0},\bar{\lambda}_{0})$ such that $F(u_{0}(s),\lambda_{0}(s))\equiv 0$ [11].
@@ 4 Concluding remarks
In conclusion, we highlight that despite the choice of discretization that respects the reflection symmetry required by (1), we get antisymmetric eigenmodes at the spurious bifurcation points on the primary solution path. Note that, the numerical solutions $u(x)$ on the primary path are all reflection symmetry-preserving. In fact, the numerical bifurcation paths even with coarser FE discretizations agree well with (2) except near the limit point. Through this work, we bring to attention that while numerical approaches are routinely used for parametrized nonlinear problems, a systematic verification of the numerical bifurcations cannot be discounted.